OpenStudy (megannicole51):

There are 12 students in a social studies class. three students will be selected to present their term projects today. In how many specific ways can three students be selected? (a) 1320 (b) 220 (c) 504 (d) 36

OpenStudy (dan815):



OpenStudy (dan815):

the question says specific ways, so i guess order matters!

OpenStudy (campbell_st):

well if there is no order to the selection its \[\frac{12 \times 11\times 10}{3 \times 2\times 1}\] the number of ways of selecting 3 people from 12 divided by the number of ways of arranging the 3 people..

OpenStudy (dan815):